# Cognitive Bias in Decision-Making with LLMs

### v2

@@ 4.3 Self-Help: Can LLMs debias their own prompts?
Mitigating patterns similar to human cognitive bias in LLMs presents two complex challenges. First, devising a specific example to illustrate a single cognitive bias is difficult, often requires long context, and it is impossible to create a generalized example that encompasses multiple biases due to their significant differences. Second, the introduction of new information can unintentionally lead to the emergence of alternative biases Teng (2013), complicating the development of examples [^1] Similar problems exist in the cognitive science literature Leung et al. (2022).. In few-shot settings, examples must be carefully crafted to be representative without introducing new biases, a process that can require extensive trial and error depending on the use-case and the number of biases involved.
Given these challenges, we explore the potential of self-help, an entirely unsupervised method where the model is tasked with rewriting prompts to mitigate cognitive bias. This approach follows a generalized process regardless of the specific bias, and offers a simple and scalable alternative to manually developing examples. In our study, we focus on one bias at a time. However, self-help can also be used iteratively to remove multiple biases. We assess the effectiveness of generating de-biased prompts by instructing the model to re-write the original question.
@@
Biased 0.002 0.000 0.002* 0.257 0.578 -0.321 0.079
Table 3: For framing and group attribution bias, we evaluate the difference in admission rate between the two (admit/reject or male/female) setups. For anchoring bias we show decision confidence in terms of normalized Euclidean distance between the general admission distribution and the (aggregated) admission distribution for individual students at different orders. We see that models show different confidence with different mitigation techniques but improve compared to the original setup. (*) indicates model failure to adhere to instructions (<1% admission or rejection ratio)
@@
This method requires no manual adaptation, but for each sample an additional forward pass is necessary. For self-help for anchoring bias, the prompts themselves can not be “de-biased” (due to the bias being induced by previous decisions). We allow the model to de-bias its own decisions based on its last prompt in the sequential procedure, which lists all student profiles and previous decisions. We ask it to change its decisions if there was a chance of bias.
@@ 5 Results
We evaluate four language models with different capabilities. We evaluate state-of-the-art commercial language models GPT-3.5-turbo and GPT-4222For group attribution and framing in GPT-4, we limit the evaluation to 400 prompts per experiment to reduce cost. These biases are not sensitive to order, so we assume the results generalize to the full data., as well as open-source large language models Llama 2 in sizes 7B and 13B.
@@ 5.1 LLMs Display Patterns Analogous to Human Cognitive Bias
In anchoring bias, we observe the existence of smaller decision confidence in the original (random order) evaluation setup, potentially attributed to the influence of previous decisions on the next decisions and unawareness of bias (Figure 3).
@@
We observe decision inconsistencies similar to human cognitive bias for framing bias and group attribution bias as shown in Table 3, where we see that all models show different behavior for admission/rejection framing and male/female group attribution. We see that GPT-4 is specifically vulnerable to patterns of framing bias where it admits 40% more students in the reject framing. Llama-2 7B is specifically vulnerable to behavior akin to human group attribution bias where the model classifies 32% fewer females as being good at math.
We do not observe a clear indication of decision patterns indicating similarities to status quo bias that is similar to human bias. Rather, we observe that for all models except GPT-4, status-quo-biased prompts are inversely biasing the model. For example, when prompting the model for the current option being option A, A is selected fewer times compared to the neutral prompt (Figure 3).
@@
Figure 3: We observe a strong primacy effect, with first options (a, b) being selected more frequently than later ones (c,d), even though all options are equally likely. Counterfactual and contrastive methods lead to failure cases that disregard options of the answer set. Self-help leads to a more balanced selection distribution. For status quo, we observe that the status quo prompting inversely biases the model to select the status quo option less frequently.
We observe that models tend to have a preference for options that are shown early in the prompt ( e.g. A or B in single-choice setup) which we see in the distribution of option selection in Figure 3, where the fraction of chosen options A or B exceeds the fraction of C plus D.
@@ 5.2 Zero-Shot Debiasing helps to mitigate Bias
In general, we see small improvements when using zero-shot prompting. For Llama models, the awareness debiasing strategy shows the best results for anchoring bias, whereas other (few-shot) methods lead to failure cases. Awareness mitigation mitigates mitigate patterns of primacy bias to a certain extent for LLama 2 and GPT-4, but does not achieve similarly good results as self-help.
@@ 5.3 Few-Shot Debiasing Can Lead to Failures
For different biases we see that few-shot prompting can lead to failure cases, e.g. driving the probability of admission/rejection to zero or one and hence undermining the ability to follow the instruction correctly for all biases, e.g. for testing for patterns of status quo bias, anchoring bias, framing or group attribution bias (Table 3), specifically for open-source LLMs. Counterfactual mitigation adds a large amount of additional context which can change the prompt drastically, lead to extreme results and loss of instruction following.
To mitigate patterns similar to human cognitive bias, giving an example often needs an explanation of the setup that leads to the bias. It can be hard to find short examples that explain the failure case sufficiently.
5.4 Models Can Debias Themselves
We see that self-help increases the decision confidence for commercial GPT models, but not for open-source Llama models (Figure 4). When allowed to change its decisions, we see that Llama models tend to change between 40-52% of their decisions, which indicates a severe amount of inconsistency in decisions between the sequential setup and the self-help setup, where all information and decisions are seen at once. We hence conclude that self-help for anchoring can only be performed by high-capacity models, or that only high-capacity models should be used to debias these prompts for lower-capacity models. For Llama models, the awareness debiasing mitigation strategy shows the best results, as contrastive and counterfactual methods either lead to low confidence or the possibility for collapse (leading to only responding with “admit” e.g. for Llama-2-7b counterfactual) (Table 4).
@@ 5.4 Models Can Debias Themselves
Table 4: When given the opportunity to change their decisions post-hoc with an overview of all student information and given an instruction to de-bias their own decisions, Llama changes their decisions frequently.
Self-help Balances Inherent Primacy Bias
Primacy bias is defined through the pattern of preference of selection for information that is first encountered. We observe in Figure 3 that the fraction of initially seen answer options (a or b) is selected more frequently compared to later options (c or d). Cognitive bias awareness prompting mitigates the issue to a small extent for Llama 2 and GPT-4, and balances the answer distribution to the desired distribution for Llama 2 7B and GPT-4. Lower capacity models like GPT-3.5-turbo have less capacity to debias themselves, but compared to other approaches that can exhibit complete failure (e.g. counterfactual prompting), self-help performs best.
Figure 4: Ratio of biased prompts that were successfully de-biased, with bias-inducing parts removed in the self-help de-biased prompt. Higher Capacity Models experience greater self-help debiasing success for prompt-induced cognitive bias.
Self-help Finds Biased Parts of the Prompt
When looking at bias which is induced by the prompt, we analyse the behavior of self-help to remove the parts of the prompt that are associated with the cognitive bias condition. We see that self-help can reduce the number of biased prompts (e.g. gender) to 0 for high-capacity models (group attribution bias - GPT-4), but fail for others (Llama). We see high debiasing performance of low capacity methods for framing bias (0% for Llama 2 13B and 1.4% for Llama 2 7B) and status quo bias, which is reduced to 6% remaining biased prompts for Llama 2 7B, 0% for Llama 2 13B. GPT-4 reduces group attribution bias elements to 0% and 2.7% for framing bias elements of the prompt. We show examples of self-help debiased prompts in Appendix Table 5. GPT-3.5 shows limited capabilities to reduce biased group attribution prompts (reduction by 8.9%), but reduces the number of biased prompts in framing and status quo to 17.2% and 8.5%.
Higher Capacity Models Experience Greater Self-help Debiasing Success
Our findings indicate fewer biased behavior of higher capacity models using self-help debiasing. These models demonstrate a notable proficiency in autonomously rewriting their input prompts to mitigate decision patterns of cognitive biases compared to lower parameter models. We specifically observe an increased number of prompts without cognitive bias-inducing words (Figure 4). Specifically, high-capacity models can reduce the bias in prompts to 0 for Group Attribution and Framing bias.
@@ 6 Conclusion
A model showing patterns similar to human cognitive bias can make inconsistent decisions, which can lead to unfair treatment in high-stakes decision-making. Our work provides a dataset of 16,800 decision questions to test for inherent, prompt-based, and sequential cognitive bias. We propose metrics to evaluate patterns of different kinds of biases and mitigation procedures. Our mitigation procedures include a novel self-debiasing technique that enables models to autonomously rewrite their own prompts, successfully removing bias inducing parts of the prompt and enabling more consistent decisions in LLMs. We observe our self-debiasing technique to be specifically successful in high-capacity models. Our method has the advantage of not requiring manually developed examples as de-biasing information to give to the model and is applicable to a variety of biases.
@@ 7 Limitations and Future Work
This work aims to encourage a protocol for continuous testing of LLMs for signs of bias that functionally resemble human cognitive bias. Our data can be used to test for LLM decision inconsistencies automatically at scale based on the final decision outcomes with small changes in the prompts. We publish our data under CC-BY NC license. The intended use of this data is to advance and facilitate the mitigation of inconsistent decisions due to cognitive bias in LLMs for high-stakes decision-making. We acknowledge the use and limitations of synthetic data as a way of analysing high-stakes decision tasks without the privacy concerns of real-world data.
@@
We discourage using LLMs independently for these high-stakes decisions, as it is important not to replicate the history of using biased automated techniques in student admissions Hutchinson and Mitchell (2019). Instead, we hope our work sheds light on the inconsistencies associated with using LLMs for high-stakes decision tasks. In future work, we aim to analyze different reasoning processes of models for their individual decisions to better assess the impact of these decisions on humans when used in human-AI collaboration setups.
@@
We examine the presence of various cognitive biases in leading commercial and open-source language models. We select a set of biases relevant to high-stakes decision-making and analyse prompts that demonstrate each bias individually. Our methodology allows for flexibility beyond single-bias testing and can accommodate multiple biases simultaneously, either through repeated applications of our technique or by modifying the prompts to target multiple biases for de-biasing. In some instances, we see multiple biases being removed (e.g. gender information in framing bias prompts (Table 5)). However, the interaction of multiple cognitive biases remains underexplored in human research, with limited research outside of analyses pertaining to specific psychological disorders (e.g. Hirsch et al. (2006); Everaert et al. (2012)), thus posing challenges in developing prompts based on human studies. Future research should focus on creating specialized testing procedures and prompts to explore the intricate dynamics between multiple cognitive biases in models, which may differ from their interplay in humans.
All experiments are run with on NVIDIA RTX A6000 (open-source models) or by querying the official APIs with a fixed random seed.
@@ A.1.1 Status Quo
In status quo bias, participants can be faced with the scenario of selecting a new car while being informed that they currently own a Honda Civic. They are then asked which car they would prefer to purchase next: (a) Hyundai Venue, (b) Honda Civic, or (c) BMW X1. This setup introduces bias by mentioning their current car situation, possibly influencing their decision towards maintaining the status quo Samuelson and Zeckhauser (1988).
@@ A.1.4 Framing
In previous work, participants were asked to choose between two treatment options for a deadly disease. Option 1 guaranteed saving 200 out of 600 people, while Option 2 offered a one-third chance of saving everyone. The study explored reactions to the same scenario framed differently, focusing on the likelihood of death rather than survival. Despite the mathematical equivalence of the outcomes in both framings, participants’ choices varied significantly, illustrating how the presentation of choices can influence decision-making Tversky and Kahneman (1981).
@@ A.1.5 Group Attribution
Biased Prompt Self-Help Debiased Prompt
@@
Table 5: Examples of self-help debiased prompts (GPT-4). In bold are key words relevant for each bias, which are successfully removed or rephrased in the self-help mitigation.
@@
SQ/Primacy Here is an example of incorrect behavior. Try to avoid this behavior. EXAMPLE: You are currently working with Student A in your lab during a summer internship, but have received more applications from students, each with different backgrounds and strengths. a) Retain Student A: Comes from a liberal arts background, … … d) Student D: Comes from an international university with a unique cultural perspective,… Your answer was: a) Here is an example of correct behavior. EXAMPLE: You are currently working with Student A in your lab during a summer internship, but have received more applications from students, each with different backgrounds and strengths. a) Retain Student A: Comes from a liberal arts background, … … d) Student D: Comes from an international university with a unique cultural perspective, … Your answer was: c) Here is an example of incorrect behavior. Try to avoid this behavior. EXAMPLE: You are currently working with Student A in your lab during a summer internship, but have received more applications from students, each with different backgrounds and strengths. a) Retain Student A: Comes from a liberal arts background, … … d) Student D: Comes from an international university with a unique cultural perspective, … Your answer was: a)
Anchoring Here are two examples being wrongly biased by anchoring bias and the correct behavior: EXAMPLE1 Wrong behavior: First Sequence: The student studied … Your decision was: Reject. The student studied … Your decision was: Reject. The student studied … Your decision was: Admit; Second Sequence: The student studied … Your decision was: Reject. The student studied … Your decision was: Admit. The student studied … Your decision was: Admit. This example shows anchoring due to inconsistency of decisions. Please try to avoid this behavior. EXAMPLE2 Correct behavior: For the same two sequences as above, your decisions were the following:First Sequence:Your decision was: Admit, Reject, Reject Second Sequence: Your decision was: Reject, Reject, Admit This example shows no inconsistency of decisions which is desirable. Here is an example two sequences being wrongly biased by anchoring bias: EXAMPLE: First Sequence: The student studied … Your decision was: Reject. The student studied… Your decision was: Reject. The student studied … Your decision was: Admit; Second Sequence: The student studied… Your decision was: Reject. The student studied … Your decision was: Admit. The student studied … Your decision was: Admit. This example shows anchoring due to inconsistency of decisions. Please try to avoid this behavior.

### v3

@@ 4.3 Self-Help: Can LLMs debias their own prompts?
Mitigating patterns similar to human cognitive bias in LLMs presents two complex challenges. First, devising a specific example to illustrate a single cognitive bias is difficult, and often requires a long context, and it is impossible to create a generalized example that encompasses multiple biases due to their significant differences. Second, the introduction of new information can unintentionally lead to the emergence of alternative biases Teng (2013), complicating the development of examples [^2] Similar problems exist in the cognitive science literature Leung et al. (2022).. In few-shot settings, examples must be carefully crafted to be representative without introducing new biases, a process that can require extensive trial and error depending on the use case and the number of biases involved.
Given these challenges, we explore the potential of selfhelp, an entirely unsupervised method where the model is tasked with rewriting prompts to mitigate cognitive bias. This approach follows a generalized process regardless of the specific bias and offers a simple and scalable alternative to manually developing examples. In our study, we focus on one bias at a time. However, selfhelp can also be used iteratively to remove multiple biases. We assess the effectiveness of generating debiased prompts by instructing the model to rewrite the original question.
@@
baseline (biased) 0.002 0.000 0.002* 0.257 0.578 -0.321 0.079
Table 3: For framing and group attribution bias, we evaluate the difference ($\Delta$) in admission rate between the two (admit/reject or male/female) setups. For anchoring bias, we show decision confidence in terms of normalized Euclidean distance $d$ between the general admission distribution and the (aggregated) admission distribution for individual students at different orders. We see that models show different indications of bias with different mitigation techniques but mostly improve compared to the original baseline (which has biased parts in the prompts). (*) indicates model failure to adhere to instructions (<1% admission or rejection ratio), where the model suddenly starts to reject or admit almost every sample.
@@
This method requires no manual adaptation, but for each sample an additional forward pass is necessary. For selfhelp for anchoring bias, the prompts themselves can not be “debiased” (due to the bias being induced by previous decisions). We allow the model to debias its own decisions based on its last prompt in the sequential procedure, which lists all student profiles and previous decisions. We ask it to change its decisions if there is a chance of bias.
@@ 5 Results
We evaluate four language models with different capabilities. We evaluate state-of-the-art commercial language models GPT-3.5-turbo and GPT-4333For group attribution and framing for GPT, we limit the evaluation to 400 prompts per experiment to reduce cost. These biases are not sensitive to order, so we assume the results generalize to the full data., as well as open-source large language models Llama 2 in sizes 7B and 13B.
@@ 5.1 LLMs Display Patterns Analogous to Human Cognitive Bias
For human-like anchoring bias, we observe the existence of small decision confidence in the original (random order) evaluation setup, potentially attributed to the influence of previous decisions on the next decisions and unawareness of bias (Figure 3).
@@
We observe decision inconsistencies similar to human cognitive bias for framing bias and group attribution bias as shown in Table 3, where we see that all models show different behavior for admission/rejection framing and male/female group attribution. We see that GPT-4 is specifically vulnerable to patterns of framing bias where it admits 40.5% more students in the reject framing. Llama-2 7B is specifically vulnerable to behavior akin to human group attribution bias where the model classifies 32.1% fewer females as being good at math.
We do not observe a clear indication of decision patterns indicating similarities to status quo bias that is similar to human bias. We observe that for all models except GPT-4, status-quo-biased prompts are inversely biasing the model. For example, when prompting the model for the status quo option being option A, A is selected fewer times (Figure 3).
@@
Figure 3: This figure shows the answer distribution for the status quo/primacy bias prompting. We observe a strong primacy effect, with first options (A, B) being selected more frequently than later ones (C, D), even though all options are equally likely. Counterfactual and contrastive methods lead to failure cases that disregard options of the answer set. Selfhelp leads to a more balanced selection distribution. For status quo biased baseline prompting, we observe that the status quo prompting inversely biases the model to select the status quo option (A) less frequently for all models except GPT-4.
We observe that models tend to have a preference for options that are shown early in the prompt (e.g. A or B in single-choice setup), akin to primacy bias, which we see in the distribution of option selection in Figure 3, where the fraction of chosen options A or B exceeds the fraction of C plus D.
@@ 5.2 Zero-Shot Debiasing helps to mitigate Bias
In general, we see small improvements when using zero-shot prompting. For Llama models, the awareness debiasing strategy shows better results for anchoring bias, whereas other (few-shot) methods lead to failure cases (Table 3). Awareness mitigation mitigates patterns of primacy bias to a certain extent (makes the distribution more uniform) for LLama 2 and GPT-4, but selfhelp leads to better results (Figure 3).
@@ 5.3 Few-Shot Debiasing Can Lead to Failures
For different biases, we see that few-shot prompting can lead to failure cases. This drives the probability of admission/rejection to zero or one and hence undermining the ability to follow the instruction correctly for all biases, e.g. for testing for patterns of status quo bias, anchoring bias, framing or group attribution bias (Table 3).
Counterfactual mitigation adds a large amount of additional context which can change the prompt drastically, lead to extreme results and loss of instruction following. To mitigate bias patterns similar to human cognitive bias, giving an example often needs an explanation of the setup that leads to bias. It can be hard to find short examples that explain the failure case sufficiently.
5.4 Models Can Remove Bias Patterns
@@
Table 4: Anchoring bias mitigation: When given the opportunity to change their decisions post-hoc with an overview of all student information and given an instruction to debias their own decisions, Llama changes their decisions too frequently.
When allowed to change their decisions for anchoring, we see that Llama models tend to change between 40-52% of their decisions (Table 4), which indicates a severe amount of inconsistency in decisions between the sequential setup and the selfhelp setup, where all information and decisions are seen at once. We hence conclude that selfhelp for anchoring can only be performed by high-capacity models, or that only high-capacity models should be used to debias these prompts for lower-capacity models (high-capacity refers to models that have a high number of parameters and extended training).
Selfhelp Balances Inherent Patterns of Primacy Bias
Primacy bias is defined through the selection preference for information that is first encountered. We observe in Figure 3 that the fraction of initially seen answer options (A or B) is selected more frequently compared to later options (C or D). Cognitive bias awareness prompting mitigates the issue to a small extent for Llama 2 7B and GPT-4. GPT-3.5-turbo has less capacity to debias itself, but compared to other approaches that can exhibit complete failure (e.g. counterfactual prompting), selfhelp performs best.
Figure 4: Ratio of biased prompts that were successfully debiased, with bias-inducing parts removed in the selfhelp debiased prompt. Higher capacity models experience greater selfhelp debiasing success for prompt-induced cognitive bias.
Selfhelp Finds Biased Parts of the Prompt
When looking at bias which is induced by the prompt, we analyze the behavior of selfhelp to remove the parts of the prompt that are associated with the cognitive bias condition. We see that selfhelp can reduce the number of biased prompts (e.g. gender) to 0 for high-capacity models (group attribution bias – GPT-4), but fail for others (Llama). We see high debiasing performance of low capacity methods for framing bias (0% for Llama 2 13B and 1.4% for Llama 2 7B) and status quo bias, which is reduced to 6% remaining biased prompts for Llama 2 7B, 0% for Llama 2 13B. Selfhelp in GPT-4 reduces group attribution bias elements to 0% and 2.7% for framing bias elements of the prompt. We show examples of selfhelp debiased prompts in Appendix Table 5. GPT-3.5 shows limited capabilities to reduce biased group attribution prompts (reduction by 8.9%), but reduces the number of biased prompts in framing and status quo to 17.2% and 8.5%.
Higher Capacity Models Experience Greater Selfhelp Debiasing Success
Our findings indicate less biased behavior of higher capacity models using selfhelp debiasing. These models demonstrate a notable proficiency in autonomously rewriting their input prompts to mitigate decision patterns of cognitive biases compared to lower parameter models. We observe an increased number of prompts without cognitive bias-inducing words (Figure 4). Specifically, high-capacity models can reduce the bias in prompts to 0 for group attribution and framing bias.
@@ 6 Conclusion
A model showing patterns similar to human cognitive bias can make inconsistent decisions, which can lead to unfair treatment in high-stakes decision-making. Our work provides a dataset of 13,465 prompts to test for inherent, prompt-based, and sequential patterns of cognitive bias in LLMs. We propose metrics to evaluate patterns of different kinds of biases and different mitigation procedures. Our mitigation procedures include a novel self-debiasing technique for patterns of cognitive bias that enables models to autonomously rewrite their own prompts, successfully removing bias-inducing parts of the prompt and enabling more consistent decisions in LLMs. We observe our self-debiasing technique to be specifically successful in high-capacity models. This method has the advantage of not requiring manually developed examples as debiasing information to give to the model and applies to a variety of biases.
@@ 7 Limitations and Future Work
Data
This work aims to encourage a protocol for continuous testing of LLMs for signs of bias that functionally resemble human cognitive bias. Our data can be used to test for LLM decision inconsistencies automatically at scale based on the final decision outcomes. We publish our data under CC-BY NC license. The intended use of this data is to advance and facilitate the mitigation of inconsistent decisions due to cognitive bias in LLMs for high-stakes decision-making. We acknowledge the use and limitations of synthetic data as a way of analysing high-stakes decision tasks without the privacy concerns of real-world data.
@@
We discourage using LLMs independently for these high-stakes decisions, as it is important not to replicate the history of using biased automated techniques in student admissions Hutchinson and Mitchell (2019). Our work gives insights on the current extent of inconsistencies. We hope our work sheds more light on the inconsistencies associated with using LLMs for high-stakes decision tasks. In future work, we aim to analyze different reasoning processes of models for their individual decisions to better assess the impact of these decisions on humans when used in human-AI collaboration.
@@
We examine the presence of patterns resembling various cognitive biases in leading commercial and open-source language models. We select a set of biases relevant to high-stakes decision-making and analyze prompts that demonstrate each bias individually. Our methodology allows for flexibility beyond single-bias testing and can accommodate multiple biases simultaneously, either through repeated applications of our technique or by modifying the prompts to target multiple biases for debiasing. We only measure if a particular bias is mitigated, but note that our framework is applicable for repeated mitigation of multiple biases. In some instances, we see multiple biases being removed (e.g. gender information in framing bias prompts (Table 5)). However, the interaction of multiple cognitive biases is still largely underexplored in human research, with only a few studies focusing on specific psychological disorders (e.g., Hirsch et al. (2006); Everaert et al. (2012)). This gap presents challenges in creating prompts informed by human studies. Future research should focus on creating specialized testing procedures and prompts to explore the intricate dynamics between multiple cognitive biases in models, which may differ from their interplay in humans.
Experiments
All experiments are run on NVIDIA RTX A6000 (open-source models) or by querying the official APIs with fixed random seed.
@@ A.1.1 Status Quo
In the original setup of status quo bias, participants are faced with the scenario of selecting a new car while being informed that they currently own a Honda Civic. They are then asked which car they would prefer to purchase next: (a) Hyundai Venue, (b) Honda Civic, or (c) BMW X1. This setup introduces bias by mentioning their current car situation, possibly influencing their decision towards maintaining the status quo Samuelson and Zeckhauser (1988).
@@ A.1.4 Framing
In previous work, participants were asked to choose between two treatment options for a deadly disease. Option 1 guaranteed to save 200 out of 600 people, while Option 2 offered a one-third chance of saving everyone. The study explored reactions to the same scenario framed differently, focusing on the likelihood of death rather than survival. Despite the mathematical equivalence of the outcomes in both framings, participants’ choices varied significantly, illustrating how the presentation of choices can influence decision-making Tversky and Kahneman (1981).
@@ A.1.5 Group Attribution
Baseline Biased Prompt Self-Help Debiased Prompt
@@
Table 5: Examples of selfhelp debiased prompts (GPT-4). In bold are key words relevant for each bias, which are successfully removed or rephrased in the selfhelp mitigation.
@@
SQ/Primacy Here is an example of incorrect behavior. Try to avoid this behavior. EXAMPLE: You [are currently working with Student A in your lab during a summer internship, but] have received [more] applications from students, each with different backgrounds and strengths. a) Retain Student A: Comes from a liberal arts background, … … d) Student D: Comes from an international university with a unique cultural perspective,… Your answer was: a) Here is an example of correct behavior. EXAMPLE: You [are currently working with Student A in your lab during a summer internship, but] have received [more] applications from students, each with different backgrounds and strengths. a) Retain Student A: Comes from a liberal arts background, … … d) Student D: Comes from an international university with a unique cultural perspective, … Your answer was: c) Here is an example of incorrect behavior. Try to avoid this behavior. EXAMPLE: You [are currently working with Student A in your lab during a summer internship, but] have received [more] applications from students, each with different backgrounds and strengths. a) Retain Student A: Comes from a liberal arts background, … … d) Student D: Comes from an international university with a unique cultural perspective, … Your answer was: a)
Anchoring Here are two examples wrongly biased by anchoring bias and the correct behavior: EXAMPLE1 Wrong behavior: First Sequence: The student studied … Your decision was: Reject. The student studied … Your decision was: Reject. The student studied … Your decision was: Admit; Second Sequence: The student studied … Your decision was: Reject. The student studied … Your decision was: Admit. The student studied … Your decision was: Admit. This example shows anchoring due to inconsistency of decisions. Please try to avoid this behavior. EXAMPLE2 Correct behavior: For the same two sequences as above, your decisions were the following:First Sequence:Your decision was: Admit, Reject, Reject Second Sequence: Your decision was: Reject, Reject, Admit This example shows no inconsistency of decisions which is desirable. Here is an example two sequences wrongly biased by anchoring bias: EXAMPLE: First Sequence: The student studied … Your decision was: Reject. The student studied… Your decision was: Reject. The student studied … Your decision was: Admit; Second Sequence: The student studied… Your decision was: Reject. The student studied … Your decision was: Admit. The student studied … Your decision was: Admit. This example shows anchoring due to inconsistency of decisions. Please try to avoid this behavior.